Given the following function: QD = 200 - 5.25P

7.1 Derive the Regular Demand Function

7.2 Derive the Inverse Demand Function

7.3 Derive the Price and Quantity which maximizes TR.

7.4 Suppose the Supply Curve is given as Qs = 25; derive the price and quantity equilibrium.

7.5 Use the price and quantity which maximizes TR and the Price and Quantity equilibrium, drive the mid-point (Arc) elasticity and interpret the result.
Answer #1

7.1. Regular demand function is Qd = 200 - (21/4)P

7.2 Inverse demand function is (21/4)P = 200 - Qd or P = 800/21 - (4/21)Qd

7.3 Total revenue is TR = PQ = (800/21 - (4/21)Q)Q. This becomes 800Q/21 - (4/21)Q^2. Place its derivative equal to 0 and see that 800/21 - (8/21)Q = 0 or Q = 100 units. At this quantity, price is P = 800/21 - 400/21 = 400/21 or 19.05

7.4. Use Qd = Qs and get 200 - 5.25P = 25. This gives P = (200 - 25)/5.25 = 100/3 or 33.33 and quantity = 25 units.

7.5. P1 = 19.05, P2 = 33.33. Q1 = 100, Q2 = 25. Use Ed = (Q2 – Q1) / [(Q2 + Q1)/2] / (P2 – P1) / [(P2 + P1)/2] and see that ed = -2.20

Q1 Q2 P1 P2 Q2-Q1 (Q2+Q1)/2 %Q P2-P1 (P1+P2)/2 %P Ed
100.00 25.00 19.05 33.33 -75.00 62.50 -120.00 14.28 26.19 54.54 -2.20